Electroplating Postlab Questions 1. Suppose that you have the opportunity to be the first in the world to come up witha your data for the number of Coulombs of charge passed and the number of moles of value of Faraday's constant -never mind we al ready know it is 96485 C/mol e. Using copper actually lost from the copper electrode in your experiment, what would you predict for the value of Faradays constant? Show your work. What is the percent error your calculated value and the accepted value of Faraday's constant? 2. How long would you have to electroplate with a current of 0.500 A to plate out 50.0-g of lead starting from Pb(NO3)h? Show your work. Now, how long would you have to electroplate with a current of 0.500 A to plate out 50.0-g of lead if your initial solution contained PbNO3)4? Notice this will be a simple calculation if you consider your answer to Question 2. 3. Suppose you are running simultaneous electroplating operations using the same current and starting solutions of CsCI, CdCl2, and InCl. Which of the operations would yield 3.0-mol of the respective metal in the shortest period of time? Show your work. 4.
Electroplating Data Sheet Quantity Value (with units) Initial mass of copper strip after cleaning Mass of copper strip after electroplating Mass of copper lost by the copper strip during electroplating Mass of watch glass and key after electroplating: 3,07 After first weighing: After second weighing:2 After third weighing: Mass of empty watch glass Final mass of key after electroplating Initial mass of key after cleaning Experimental mass of copper deposited during electroplating Number of Coulombs of charge passed during electroplating 30 2 9. a 212 (integration of Current vs. Time plot) Number of moles of electrons passed during electroplating Number of moles of copper expected to be electroplated based on the moles of electrons Number of moles of copper actually electroplated on key Percent yield of copper based on the electroplating of the key Number of moles of copper expected to be lost from copper strip Number of moles of copper actually lost from the copper strip Percent of actual moles of copper lost by copper strip to moies expected to be lost by copper strip
Answer #1

2)

in Pb(NO3)2 the Pb is in +2 oxidation state

so by applying the current, the following reaction will take places for electroplating

Pb2+ + 2e- ------> Pb

for electroplating 50 g of lead

moles of Pb = 50 / mass of Pb = 50 / 207.2 = 0.241 mol

since one mole of Pb is formed by accepting 2 electrons

so moles of electrons = 2 * 0.241 = 0.482 mol

since 1 mol of electron contain 96485 C of charge

so total charge transfer (dQ) = 96485 * 0.482 = 46505.8 C

current = dQ / dT

dT = dQ/current = 46505.8 / 0.5 = 93011.5 second

so time = 93011.5 / 60 * 60 = 25.84 hours

so for electroplating 50 g of Pb at 0.5 A in Pb(NO3)2 solution the electroplating should be done for 25.84 hour

3)

in the case of Pb(NO3)4 the Pb is in +4 oxidation state

so by applying current the following reaction take place,

Pb4+ + 4e- -----> Pb

so

for electroplating 50 g of lead

moles of Pb = 50 / mass of Pb = 50 / 207.2 = 0.241 mol

since here Pb is formed by accepting 4 electrons

so moles of electrons = 4 * 0.241 = 0.964 mol

since 1 mol of electron contain 96485 C of charge

total charge transfer (dQ) = 96485 * 0.964 = 93011.5 C

current = dQ / dT

dT = dQ / current

dT = 93011.5 / 0.5

dT = 186023.1 second

dT = 51.67 hour

so for electroplating 50 g of Pb at 0.5 A in Pb(NO3)4 solution the electroplating should be done for 51.67 hour

so time is double compare to time in 2 question because here the no of electrons transfer are double compered to 2 question.